next door neighbors Bob and Jim use hoses from both houses to fill Bob's swimming pool. they know that it takes 18 hrs. using both hoses.they also know that bob's hose,used alone takes 20% less time than jims hose alone how much time is required to fill the pool by each hose alone?
Answer by mananth(16949)

Let Jim's hose take x hours
it does 1/x of job in 1 hour
Bob's hose will take 0.8x hours ( 20 percent less)
it does 1/0.8x of the job in 1 hour
together they take 18 hours
so they do 1/18 of the job in 1 hour
1/x + 1/0.8x = 1/18
1/x + 1.25/x = 1/18
multiply by 18x
18+18*1.25 = x
x= 40.5 hours Jim's hose
Bobs hose will tale 0.8*40.5 = 32.4 hours